OpenStudy (anonymous):

Find the x and the y intercepts for f(x)=-x^2+2x+15 The x- coordinates of the intercept are x= The y- intercept is (0, )

OpenStudy (anonymous):

when x=0, f(x)=-0^2+2(0)+15- 15 y-intercept = 15 when f(x)=0 0=-x^2+2x+15 0=(x+3)(x-5) x= -3, 5